# Inverse k-visibility for RSSI-based Indoor Geometric Mapping

### v1

@@ VI-A Evaluation Metrics
Prior to any metric evaluation, the RSSI-generated map must be aligned and superposed along the ground-truth lidar map, both discretized by an aligned occupancy grid map. Each grid element on the grid map represents a small-element subset of the entire $M\times N$ map which is used for element-wise comparison. Here, we will present the evaluation metrics used in measuring the experimental accuracy after alignment.
k-value accuracy Percentage: We use the wall prediction model proposed by Fafoutis et al. [53], as discussed further in Sec. V-A to obtain the $k$-values from the acquired RSSI values which uses a $K$-Means approach. The RSSI-generated map is then estimated by Alg. 1. From the ground-truth map, we then construct expected $k$-values based on the positioning of the routers and the wall obstacles.
Ultimately, the $k\text{-}value_{accuracy}$ is an element-by-element comparison of the estimate $k$-values from the estimate RSSI-generated map and the ground-truth $k$-values from the ground-truth map. If the estimate $k$-value is the same as the expected on the ground truth, it is deemed predicted True and,
@@
IOU Score: A mask $Z$ on each coordinate $(i,j)\in M\times N$ first needs to be applied on the RSSI-estimate $RSSIE$ map to exclude any estimates that are neither considered free-space or a wall:
$Z(i,j)=\begin{cases}1&\text{if }RSSIE(i,j)\neq 127\\ &\quad\text{and }GT(i,j)\neq 127\\ 0&\text{otherwise}\end{cases}$ (8)
where $GT$ is ground-truth map of $M\times N$. Then, the mask $Z$ is applied to filter out the uncertain parts by an intersection,
$I(i,j)=\begin{cases}1&\text{if }RSSIE(i,j)\neq 0\\ &\quad\text{and }GT(i,j)\neq 0\\ &\quad\text{and }Z(i,j)=1\\ 0&\text{otherwise}\end{cases}$ (9)
Next, account for the entire map by a union $U$ where the element is certain (either free-space or wall),
$U(i,j)=\begin{cases}1&\text{if }(RSSIE(i,j)\neq 0\\ &\quad\text{or }GT(i,j)\neq 0)\\ &\quad\text{and }Z(i,j)=1\\ 0&\text{otherwise}\end{cases}$ (10)
The IOU Score $IOU_{score}$ is then calculated by,
$IOU_{score}=\frac{\sum_{i,j}I(i,j)}{\sum_{i,j}U(i,j)}$ (11)
MSE Score: First, the difference $D_{i,j}$ is taken element-wise between $RSSIE$ and $GT$ and then squared $DS$,
$D(i,j)=RSSIE(i,j)-GT(i,j)$ (12)
$DS(i,j)=(D(i,j))^{2}=(RSSIE(i,j)-GT(i,j))^{2}$ (13)
Next, MSE Score $MSE_{score}$ is the mean of the square differences $DS$, and is calculated by,
@@
$=\frac{1}{N\times M}\sum_{i,j}(RSSIE(i,j)-GT(i,j))^{2}$
@@
VI-B Quantitative Results
Experiments 1 and 2 from SfW[5] results are extended for the single-router case to include comparison by mapping area, number of data points, $k$-value prediction True/False setting, $k$-value accuracy percentage, IOU and MSE scores. Further analysis was conducted to evaluate our proposed method, as shown in Table. I. Experiments 5-6 show improved performance with fewer data points. This is because their experiments utilize two routers instead of one. With two routers, more data is collected simultaneously, and the most optimized data is selected from various cases. In contrast, using a single router provides only one piece of information at a time causing less efficient data collection.
@@ VI-B Quantitative Results
In the first experiment, a total of 3159 data points were collected and showed 84.93% $k$-value prediction accuracy. Here, the Intersection Over Union (IOU) score was calculated at 0.8531 comparing each corresponding pixel from the ground truth and the RSSI-generated map. The Mean Squared Error (MSE) score was 1.7770 ${px}^{2}$ showing high similarity with the ground truth.
@@
Similarly, the second experiment showed 82.79% accuracy rate out of 4713 data points. The area of the map was more than twice of the first experiment. The IOU score and the MSE score was 0.9028 ${px}^{2}$ and 17.4238 ${px}^{2}$ respectively. Further analysis for the first and second experiments can be found in the prior SfW paper [5].
@@
TABLE I: Accuracy and map similarity comparison between Lidar (Ground Truth) and RSSI generated map. For brevity, $\uparrow$ means a result that is larger is better, and $\downarrow$ means a result that is lower is better.
Exp.1 Exp.2 Exp.3 Exp.4 Exp.5 Exp.6 Area of Map 15.7 m2 34.2 m2 42.5 m2 51.5 m2 42.5 m2 51.5 m2 #Router 1 1 1 1 2 2 #Data points 3159 4713 6001 1803 492 393 $k$-value Prediction True 2683 3902 5120 1698 471 379 $k$-value Prediction False 476 811 881 105 21 14 $k$-value Accuracy % $\uparrow$ 84.93 82.79 85.32 94.18 95.73 96.44 IOU Score $\uparrow$ 0.8531 0.9028 0.8321 0.9349 0.8364 0.9627 MSE Score $\downarrow$ 1.7770 17.4238 9.0783 12.1403 1.6163 13.5331
VI-C Single-Router to Multi-Router Evaluations
@@
Experiments 3-6 test the proposed algorithm for multi-router cases, where it is expected that additional router sensors would greatly improve the accuracy of the map as the WiFi signal strength is superposed and strengthened. We will start with details on single-router, then, introduce more.
@@
With the multi-router case for the fifth experiment in Fig. 8-(b), the $k$-value accuracy greatly improves going from 85.32% to 95.73% with significantly fewer data points. The second router is leveraged in the $k$ region where prior single-router cases struggled. As our algorithm makes each router to only prioritize the closest and strongest signal, the $k=0$ regime is effectively identified with minor false prediction. Details on the multi-router case for the third experiment are stated as Exp. 5 in Table I.
@@
With the multi-router case for the sixth experiment in Fig 8-(d), again, the RSSI map accuracy improves, this time going from 94.18% to 96.44%. The increased improvement is not as great as the third experiment, only with a 2% increase but this is due to the larger area of the fourth experiment. In comparing the single-router with the multi-router case, it is clear that in regions where the single-router map estimate fails, the multi-router tracks more accurately the free space area, whilst circumventing the false positives of walls. Exp. 6 accurately differentiates the free space/potential walls which were poorly detected in the $k=2$ region from the fourth experiment. Details on the multi-router case for the third experiment are stated as Exp. 6 in Table I.
VI-D Experimental Discussion Between Trajectories
@@
While RSSI data points ranging between 393 to 6001 were collected from four experiments, all experiments have shown good estimates of the $k$ values above 82%. In all four experiments, the algorithm shows a robust prediction capability of the RSSI signals across various environments in its high $k$-value accuracy percentage, its near 1 IOU score, and its relatively low MSE score, shown in Table I.
@@
The $k$-value accuracy percentage was optimized higher in experiments 4-6 in larger map areas compared to the prior experiments, by the positioning of the routers and the careful traversing of the robot trajectory to optimize RSSI signal strength from the routers. This implies that methodically planning the trajectory beforehand in the map estimate is essential.
@@
Next, the IOU score, which measures the overlap between predicted free space and the ground truth, showed little variability for all experiments ranging from 0.8-0.96, where a near 1 IOU Score indicates a good overlap between the ground truth and our algorithms.
@@
Though the $k-value$ accuracy percentage and IOU score showed great accuracy and robustness in the map estimate over all experiments, the MSE score had significant variability. Exp. 2-4 and 6 shown in Table I had larger MSE scores, but given the large area spaces, these MSE scores are within reason and alleviated when going from single-router to multi-router cases, as shown by the significant drop in MSE score of single-router Exp. 3 with MSE score of 9.0783 ${px}^{2}$ to multi-router Exp. 5 with MSE score of 1.6163 ${px}^{2}$.
Future work and Conclusions

### v2

@@ VI-A Evaluation Metrics
Prior to any metric evaluation, the RSSI-generated map must be aligned and superposed along the ground-truth LiDAR map, both discretized by an aligned occupancy grid map. Each grid element on the grid map represents a small-element subset of the entire $M\times N$ map, which is used for element-wise comparison. Here, we will present the evaluation metrics used in measuring the experimental accuracy after alignment.
k-value accuracy Percentage: We use the wall prediction model proposed by Fafoutis et al. [fafoutis2015rssi], as discussed further in Sec. V-A to obtain the $k$-values from the acquired RSSI values which uses a $K$-Means approach. The RSSI-generated map is then estimated by Alg. 2. From the ground-truth map, we then construct expected $k$-values based on the positioning of the routers and the wall obstacles.
Ultimately, the $k\text{-}value_{accuracy}$ is an element-by-element comparison of the estimate $k$-values from the estimate RSSI-generated map and the ground-truth $k$-values from the ground-truth map. If the estimated $k$-value is the same as the expected on the ground truth, it is deemed predicted True and,
@@
IOU Score: A mask $Z$ on each coordinate $(i,j)\in M\times N$ first needs to be applied on the RSSI-estimate $RSSIE$ map to exclude any estimates that are neither considered free-space nor a wall:
$Z(i,j)=\begin{cases}1&\text{if }RSSIE(i,j)\neq 127\\ &\quad\text{and }GT(i,j)\neq 127\\ 0&\text{otherwise}\end{cases}.$ (10)
where $GT$ is ground-truth map of $M\times N$. Then, the mask $Z$ is applied to filter out the uncertain parts by an intersection:
$I(i,j)=\begin{cases}1&\text{if }RSSIE(i,j)\neq 0\text{ and }GT(i,j)\neq 0\\ &\quad\text{and }Z(i,j)=1\\ 0&\text{otherwise}\end{cases}.$ (11)
Next, we account for the entire map by a union $U$ where the element is certain (free-space or wall):
$U(i,j)=\begin{cases}1&\text{if }(RSSIE(i,j)\neq 0\text{ or }GT(i,j)\neq 0)\\ &\quad\text{and }Z(i,j)=1\\ 0&\text{otherwise}\end{cases}.$ (12)
The Intersection Over Union (IOU) Score, $IOU_{score}$, is then calculated by
$IOU_{score}=\frac{\sum_{i,j}I(i,j)}{\sum_{i,j}U(i,j)}.$ (13)
MSE Score: First, the difference $D_{i,j}$ is taken element-wise between $RSSIE$ and $GT$ and then squared $DS$:
D(i,j)=RSSIE(i,j)GT(i,j),D(i,j)=RSSIE(i,j)-GT(i,j), (14)
$DS(i,j)=(D(i,j))^{2}=(RSSIE(i,j)-GT(i,j))^{2}.$ (15)
Next, the Mean Squared Error (MSE) Score, $MSE_{score}$, the mean of the square differences $DS$, is calculated by
@@
$=\frac{1}{N\times M}\sum_{i,j}(RSSIE(i,j)-GT(i,j))^{2}.$
@@
Also, the parameters referred to are the propagation model parameters, and these were first addressed in Sec. V-A and Sec. V-C. The bounds on the RSSI parameter $t_{k}$ (in Eq. (5) of our work) in the real-world experiments are found using the centroids provided by the K-Means algorithm on the aligned RSSI-generated map and superposed along the ground-truth LiDAR map for scaling consistency. This clustering provides centroids that serve as thresholds to differentiate different levels of visibility ($k$-values), ensuring consistent scale alignment with the LiDAR-based ground truth. The number of clusters used in the experiment is three as $k$-value of 0, 1, and 2 were defined.
Figure 10: Real-world experimental results for single-router and multi-router SfW mapping. (a) and (b) are for the same environment, but with one and two routers, respectively. The same applies to (c) and (d). Our proposed method with RSSI-based mapping detects most of the free space and potential walls with $k\geq 1$. The multi-router case has improved accuracy in estimating its map. Note: For each pair from (a)-(d), the left shows the ground truth obtained from LiDAR, while the right shows the SfW mapping.
In Eq. (6), the probabilistic certainty $\mu_{j}$ of the $j$-th cell being a wall in our LOS point along the robot’s trajectory is handled within the algorithm rather than shown in each instance, given the large number of trajectories and overlapping segments encountered. Similarly, Eq. (7), where the combined probability of the cell by multi-ray intersection updating $\mu$ is done in the algorithm to reflect superposed probabilities from intersecting rays, including $\mu_{1}$ and $\mu_{2}$, without needing to detail every recalculation.
For real-world adjustments, the $\sigma$ parameter associated with free space probabilities along each pixel’s LOS path was specifically tuned. This $\sigma$ value, assumed to follow a Gaussian distribution based on RSSI fluctuations near each pixel, was refined to closely align the RSSI-derived map with the ground truth obtained from LiDAR. Different $\sigma$ values were tested based on data volume, with larger datasets providing more overlapping trajectories. With more overlap, incremental updates to $\mu$ values can be smaller, balancing the density of trajectory data with accurate map representation. This introduces a trade-off: dense data allows smaller $\sigma$ adjustments, while sparse data benefits from slightly larger adjustments to maintain consistency across the map. In Fig. 10, the model was tested with a $\sigma$ pixel value of 10.
VI-B Discussion: Ignoring $k$-visibility
Figure 11: Real-world results of the comparison of mapping results between LOS alone (left top and bottom) and the inverse $k$-visibility algorithm (right top and bottom). LOS alone is insufficient when wanting to also have a measure of spatial reliability on the mapping prediction.
Using Line-of-Sight (LOS) alone will not provide enough information to give a very fine and detailed prediction of the entire map. This is treating all $k$-values as $k=0$, which would result in mapping the entire space as free. In this case, only WiFi signals are used, resembling applications such as a person walking in an environment to map it using only a WiFi receiver, or a robot with only a WiFi receiver. The inverse $k$-visibility algorithm provides a more sophisticated approach by distinguishing between different visibility values. In using a set of different $k$-values, the detail of the mapping prediction is provided, which would be a strong metric or basis to use in knowing the trustworthiness of the prediction. If there were no varying $k$-value with all $k=0$, then it would be hard to determine which parts of the predicted map are accurate enough, especially in robotic applications where mapping accuracy is essential. The comparison of doing mapping with LOS alone compared with the algorithm is provided in Fig. 11. All in all, this will ensure a more accurate representation of the space, particularly in environments where walls and obstacles exist.
VI-C Quantitative Results
For all experiments, CPU 12th Gen Intel(R) Core(TM) i7-12650H was used to evaluate different methods. The experiments were performed in an indoor environment with walls and closed layouts, chosen specifically to ensure that direct-path signals dominate the received measurements. Note that no existing WiFi-only method outputs a directly comparable occupancy map under the same sensing assumptions and evaluation protocol. Prior works are primarily aimed at RSSI heat map generation [zou2020adversarial] or robot/access-point localization rather than explicit occupancy mapping [ferris2007wifi, arun2022p2slam]. Approaches that reconstruct occupancy from wireless measurements employ different sensing hardware or experimental setups, such as directional antennas [gonzalez2013cooperative]. Because these differences affect both the available information and the evaluation conditions, we report internal metrics (IoU, MSE) computed on aligned grids against LiDAR ground truth within our test environments rather than a side-by-side baseline. Experiments include comparison by mapping area, number of data points, $k$-value prediction True/False setting, $k$-value accuracy percentage, IOU, and MSE scores. Further analysis was conducted to evaluate our proposed method, as shown in Table I. Experiments 5-6 show improved performance with fewer data points. This is because their experiments utilize two routers instead of one. With two routers, more data is collected simultaneously, and the most optimized data is selected from various cases. In contrast, using a single router provides only one piece of information at a time, causing less efficient data collection.
@@ VI-C Quantitative Results
In the first experiment, a total of 3159 data points were collected and showed 84.93% $k$-value prediction accuracy. Here, the IOU score was calculated at 0.8531, comparing each corresponding pixel from the ground truth and the RSSI-generated map. The MSE score was 1.7770 ${px}^{2}$, showing high similarity with the ground truth.
@@
Similarly, the second experiment showed an 82.79% accuracy rate out of 4713 data points. The area of the map was more than twice of the first experiment. The IOU score and the MSE score were 0.9028 ${px}^{2}$ and 17.4238 ${px}^{2}$ respectively.
@@
TABLE I: Accuracy and map similarity comparison between LiDAR (Ground Truth) and RSSI generated map. For brevity, $\uparrow$ means a larger result is better, and $\downarrow$ means a lower result is better.
Exp.1 Exp.2 Exp.3 Exp.4 Exp.5 Exp.6
Area of Map 15.7 m2 34.2 m2 42.5 m2 51.5 m2 42.5 m2 51.5 m2
#Router 1 1 1 1 2 2
#Data points 3159 4713 6001 1803 492 393
Total computation time 32.94 s 34.05 s 46.18 s 30.75 s 3.04 s 2.88 s
Time per data 5.21 ms 3.61 ms 3.85 ms 8.53 ms 3.09 ms 3.66 ms
$k$-value Prediction True 2683 3902 5120 1698 471 379
$k$-value Prediction False 476 811 881 105 21 14
$k$-value Accuracy % $\uparrow$ 84.93 82.79 85.32 94.18 95.73 96.44
IOU Score $\uparrow$ 0.8531 0.9028 0.8321 0.9349 0.8364 0.9627
MSE Score $\downarrow$ 1.7770 17.4238 9.0783 12.1403 1.6163 13.5331
VI-D Single-Router to Multi-Router Evaluations
@@
Experiments 3-6 test the proposed algorithm for multi-router cases, where it is expected that additional router sensors would greatly improve the accuracy of the map as the WiFi signal strength is superposed and strengthened. We will start with details on single-router, then introduce more.
@@
With the multi-router case for the fifth experiment in Fig. 10-(b), the $k$-value accuracy greatly improves, going from 85.32% to 95.73% with significantly fewer data points. The second router is leveraged in the $k$ region where prior single-router cases struggled. As our algorithm makes each router to only prioritize the closest and strongest signal, the $k=0$ regime is effectively identified with minor false prediction. Details on the multi-router case for the third experiment are stated as experiment 5 in Table I.
@@
With the multi-router case for the sixth experiment in Fig 10-(d), again, the RSSI map accuracy improves, this time going from 94.18% to 96.44%. The increased improvement is not as great as the third experiment, only with a 2% increase, but this is due to the larger area of the fourth experiment. In comparing the single-router with the multi-router case, it is clear that in regions where the single-router map estimate fails, the multi-router tracks more accurately the free space area, whilst circumventing the false positives of walls. Experiment 6 accurately differentiates the free space/potential walls, which were poorly detected in the $k=2$ region, from the fourth experiment. From the evaluation metric, the computational time per RSSI was calculated to be between 3ms to 8ms. The exact number can vary depending on the hardware capacity. However, it shows that the algorithm can process the RSSI in real-time. Details on the multi-router case for the third experiment are stated as experiment 6 in Table I.
In our implementation, the “focused router” is chosen as the one with the strongest RSSI, and updates are applied along that router’s ray only. While this simplifies updates and improves $k=0$ detection, it can leave bridging regions between routers unobserved when the trajectory remains within one router’s dominance zone if the trajectory does not cross-link the zones. We recommend targeting that band with active exploration (frontier/information-gain moves and short cross-connecting passes). See Sec. VIII for details and references.
VI-E Discussion
@@
While RSSI data points ranging between 393 to 6001 were collected from four experiments, all experiments have shown good estimates of the $k$ values above 82%. These four experiments show a robust prediction capability of the RSSI signals across various environments in its high $k$-value accuracy percentage, its near 1 IOU score, and its relatively low MSE score, shown in Table I.
@@
The $k$-value accuracy percentage was optimized higher in experiments 4-6 in larger map areas compared to the prior experiments, by the positioning of the routers and the careful traversing of the robot trajectory to optimize RSSI signal strength from the routers. This implies that methodically planning the trajectory in the map estimate is essential.
@@
The IOU score, which measures the overlap between predicted free space and the ground truth, showed little variability for all experiments, ranging from 0.8-0.96, where a near 1 IOU Score indicates a good overlap between the ground truth and our algorithms.
@@
Though the kvaluek-value accuracy percentage and IOU score showed great accuracy and robustness in the map estimate over all experiments, the MSE score had significant variability. Experiments 2-4 and 6 shown in Table I had larger MSE scores, but given the large area spaces, these MSE scores are within reason and alleviated when going from single-router to multi-router cases, as shown by the significant drop in MSE score of single-router experiment 3 with MSE score of 9.0783 ${px}^{2}$ to multi-router experiment 5 with MSE score of 1.6163 ${px}^{2}$.
VI-F Sensitivity analysis
In the same simulation setup as Fig. 8, we evaluated robustness to two uncertainties: (i) router-coordinate noise, modeled as zero-mean Gaussian perturbations with standard deviation $\sigma\in\{0,1,2,5,10,20\}$ px (1 px $\approx$ 0.025 m) and (ii) odometry drift, modeled as zero-mean Gaussian increments that accumulate over time with per-step standard deviation $\sigma\in\{0,0.1,0.2,0.5,1.0,2.0\}$ px. The smaller per-step values for odometry reflect realistic compounding drift in practice. For router coordinate perturbations, performance remained essentially stable (Table II), indicating overall robustness to modest placement errors. However, this robustness is conditional on line-of-sight (LOS). When router trajectory rays remain predominantly clear (e.g., centrally placed routers with broad coverage), the geometric cues are informative and performance is stable. In contrast, unfavorable placements (e.g., corner locations or frequent occlusions) reduce the number of useful sightlines, which degrades the spatial informativeness of the measurements and thus tends to decrease IoU and increase MSE. On the other hand, odometry with accumulating noise degrades map quality sharply: small per-step perturbations compound over time, and for noise standard deviation $\geq 0.5$, the drift overwhelms wall-based constraints, leading to failure to complete the map. These results underscore the need for re-localization or sensor fusion to bound long-term error (Table III). Visualizations for router-coordinate noise and odometry noise can be found in Fig. 12 and Fig. 13.
TABLE II: Router coordinate sensitivity (Gaussian perturbation of perceived router locations).
   Noise Std (px)    IoU $\uparrow$    MSE $\downarrow$
   0    0.9556    7.45
   1    0.9554    7.49
   2    0.9552    7.48
   5    0.9549    7.33
   10    0.9542    7.27
   20    0.9538    7.91
TABLE III: Odometry sensitivity under cumulative drift. Fail indicates the run did not complete successfully (drift caused divergence).
   Noise Std (px)    IoU $\uparrow$    MSE $\downarrow$
   0    0.9556    7.45
   0.1    0.9468    8.75
   0.2    0.9085    22.87
   0.5    Fail    Fail
   1.0    Fail    Fail
   2.0    Fail    Fail
Figure 12: Sensitivity of SfW mapping to router coordinate perturbations. Top row (left to right): $\sigma=0$, 1, 2 px; bottom row: $\sigma=5$, 10, 20 px. The occupancy maps remain stable with only minor degradation observed at higher noise levels. The green circles show the unperturbed positions of the routers, and the red circles show the perturbed positions of the routers.
Figure 13: Sensitivity of SfW mapping to odometry drift. Top row: $\sigma=0$, 0.1 px; bottom row: $\sigma=0.2$, 0.5 px. The green trajectories show the unperturbed trajectory of the robot, and the red trajectories show the perturbed trajectories of the robot. Higher drift causes divergence and leads to failure with cases observed in Table III.
VII Discussion and Limitations
In this work, we intentionally simplified the RSSI propagation model, assuming that received signal strength is primarily determined by the direct path through thick, attenuating walls, to validate the core concept of inverse k-visibility in a proof-of-concept setting. All experiments were carried out in a controlled indoor environment with thick interior walls and minimal open corridors, such that non-line-of-sight reflections, diffractions, and scattering were naturally dampened. Under these conditions, variations in RSSI could be attributed primarily to the number of intervening walls, making the direct-path approximation reasonable. This also enables the isolation and demonstration of the feasibility of RSSI-based geometric mapping without interference from multipath effects.
However, we fully acknowledge that this modeling choice limits robustness in more general indoor scenarios. In real-world office or home environments with open-plan areas, long hallways, reflective ceilings, and dynamic agents (i.e., moving people) can dramatically alter RSSI measurements. Our current sliding window filtering is insufficient to handle these effects, and the algorithm as presented will struggle to detect and accommodate rapidly changing signal paths. Addressing these challenges will require advanced signal processing strategies such as adaptive filtering, machine learning based denoising, or incorporating additional channel metrics (i.e., multi-band analysis) to distinguish direct-path attenuation from multipath fluctuations.
Another limitation is that our implementation is strictly 2D and planar, with both routers and trajectories confined to a ground-level grid. While this choice simplified initial validation, it does not capture ceiling-mounted access points or multi-level structures. Extending the framework to full 3D environments where vertical signal paths and signal transmission between floors become significant will be an essential next step for broader application.
We also observed that model performance degrades once the true number of walls exceeds two (i.e., $k$ $\geq 3$). In these higher $k$ regimes, our algorithm becomes less reliable, and the algorithm’s ability to accurately delineate adjacent $k$-regions diminishes. This limitation again comes from our coarse RSSI-to-$k$-mapping and the lack of explicit multipath modeling.
VIII Future work and Conclusions